OpenStudy (anonymous):

Calculus AB question. Water is flowing into a cone-shaped tank at rate of 5 cubic inches per second. If the cone has an altitude of 4 inches and a base radius of 3 inches, how fast is the water level rising when the water is 2 inches deep? Be sure to include appropriate units in your answer.

OpenStudy (dumbcow):

you need volume of a cone \[V = \frac{1}{3}\pi r^{2} h\] and ratio of height to radius \[\frac{h}{r} = \frac{4}{3}\] since they want rate h is increasing, get V in terms of only "h" \[r = \frac{3h}{4}\] \[V = \frac{3}{16} \pi h^{3}\]

OpenStudy (dumbcow):

from there, - differentiate with respect to "t" (implicit) - plug in dV/dt = 5 - solve for dh/dt - plug in h=2

OpenStudy (anonymous):

How did you go from 4/3 to 3h/4?

OpenStudy (dumbcow):

i solved for "r"

OpenStudy (dumbcow):

its just a proportion, i could have also said r/h = 3/4

OpenStudy (anonymous):

How did u go from r=3h/4 to V = 3/16 pi h^2? When you plugged r in, how did u get V = 3/16 pi h^2?

OpenStudy (dumbcow):

after you square "r" you just simplify...i figured you could simplify that without seeing every step (3h/4)^2 = 9h^2/16 * 1/3 = 3h^2/16 * h = 3h^3/16

OpenStudy (anonymous):

How do you do "differentiate with respect to "t" (implicit)"?